OpenStudy (anonymous):

WILL GIVE MEDAL!!! In a bag containing black and white marbles,3/8 of the marbles are black. Which could be the numbers of black and white marbles in the bag? A. 3 black and 8 white B. 12 black and 20 white C. 20 black and 12 white D. 8 black and 3 white

OpenStudy (bibby):

find the ratio where 3:8 represents the number of blacks to whites

OpenStudy (anonymous):

how?

OpenStudy (bibby):

what I'd do is set them all as fractions of black to white. so for example b = \(\frac{12}{20}\) and then just simplify it \(\large \frac{12}{20}=\frac{3}{5}\)